OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

T+10=C C-3=2*(T-3) C=coach T=Tim Now solve for the system of equations. Can you do it now?

OpenStudy (anonymous):

i get 30 for the answer but its not right

OpenStudy (anonymous):

thanks for your help